# A Survey and Benchmark of Automatic Surface Reconstruction from Point Clouds

### v3

@@ Abstract
We present a comprehensive survey and benchmark of both traditional and learning-based methods for surface reconstruction from point clouds. This task is particularly challenging for real-world acquisitions due to factors like noise, outliers, non-uniform sampling, and missing data. Traditional approaches often simplify the problem by imposing handcrafted priors on either the input point clouds or the resulting surface, a process that can necessitate tedious hyperparameter tuning. Conversely, deep learning models have the capability to directly learn the properties of input point clouds and desired surfaces from data.
We study the influence of these handcrafted and learned priors on the precision and robustness of surface reconstruction techniques. We evaluate various time-tested and contemporary methods in a standardized manner. When both trained and evaluated on point clouds with identical characteristics, the learning-based models consistently produce superior surfaces compared to their traditional counterparts—even in scenarios involving novel shape categories. However, traditional methods demonstrate greater resilience to the diverse array of point cloud anomalies commonly found in real-world 3D acquisitions. For the benefit of the research community, we make our code and datasets available, inviting further enhancements to learning-based surface reconstruction. This can be accessed at https://github.com/raphaelsulzer/dsr-benchmark.
@@ 3.3 Reconstruction
qtree/.style=for tree=parent anchor=south, child anchor=north,align=center,inner sep=0pt,l=1.2cm
@@ 4.3 Rendering-based reconstruction
Our paper specifically addresses surface reconstruction methods from 3D point clouds. However, a new emerging field is the reconstruction of surfaces directly from images using rendering techniques from computer graphics. Originally, so-called novel view synthesis methods [77, 78] were specifically designed for generating new views of an object or scene starting from a set of given images with their accurate pose. These methods present many advantages, including the ability to generate photo-realistic content, to deal with complex lighting and transparencies [79], and the fact that they do not rely on complex sensors. Recently, more and more rendering-based methods have been designed specifically for surface reconstruction [80, 81, 82]. They can be divided into the two categories, neural radiance fields and 3D Gaussian splatting and are closely related to the two corresponding, more general, principles from computer graphics, namely ray tracing and rasterization.
4.3.1 NeRF approaches
NeRF encode a radiance field in a neural network. The network maps from each 3D location and 2D viewing direction to a volume density and directional emitted color [77]. The color of any ray passing through the scene is rendered using principles from classical volume rendering [83]. Rendering rays from the given input images and poses allows to train the network weights, and to generate novel views from the trained network. Similar to the NIF methods that encode signed distance, or occupancy fields presented above, the volume density of neural radiance fields can be used to infer an object’s surface. However, the trivial solution of extracting the surface as a level-set of the volume density of a NeRF does not lead to a satisfying surface [80]. A common approach is to force the volume density field to change abruptly, indicating the presence of surface [81].
The benefit of ray tracing-based approaches is that the two-step approach of point cloud generation from images and subsequent surface reconstruction from the point clouds is solved in one end-to-end optimization. The downside is that the methods are often slow due to frequent querying of the radiance field during optimization.
4.3.2 3D Gaussian splatting approaches
A more efficient rendering technique compared to ray tracing is rasterization. 3D Gaussian Splatting [84] is the first method that makes use of a differentiable rasterization to generate novel views of an object or scene. The method takes a sparse point cloud, typically from SfM as input. The points are then used to initialize 3D Gaussians which are rendered using rasterization. The position and density of the 3D Gaussians is optimized using an image-based loss between input and rendered views.
The method is significantly faster for novel view synthesis compared to NeRF. However, for surface reconstruction, 3D Gaussians still need to be processed in a seperate step, e.g., using Poisson surface reconstruction [82], leading to a two step solution, similar to a traditional MVS pipeline.

### v4

@@ Abstract
We present a comprehensive survey and benchmark of both traditional and learning-based methods for surface reconstruction from point clouds. This task is particularly challenging for real-world acquisitions due to factors such as noise, outliers, non-uniform sampling, and missing data. Traditional approaches often simplify the problem by imposing handcrafted priors on either the input point clouds or the resulting surface, a process that can require tedious hyperparameter tuning. In contrast, deep learning models have the capability to directly learn the properties of input point clouds and desired surfaces from data. We study the influence of handcrafted and learned priors on the precision and robustness of surface reconstruction techniques. We evaluate various time-tested and contemporary methods in a standardized manner. When both trained and evaluated on point clouds with identical characteristics, the learning-based models consistently produce higher-quality surfaces compared to their traditional counterparts—even in scenarios involving novel shape categories. However, traditional methods demonstrate greater resilience to the diverse anomalies commonly found in real-world 3D acquisitions. For the benefit of the research community, we make our code and datasets available, inviting further enhancements to learning-based surface reconstruction. This can be accessed at https://github.com/raphaelsulzer/dsr-benchmark.
@@ 3.3 Reconstruction
[…]
@@ 4.3 Rendering-based reconstruction
Our paper specifically addresses surface reconstruction methods from 3D point clouds. However, a new emerging field is the reconstruction of surfaces directly from images using rendering techniques from computer graphics. Originally, so-called novel view synthesis methods [77, 78] were specifically designed for generating new views of an object or scene starting from a set of given images with their accurate pose. These methods present many advantages, including the ability to generate photo-realistic content, to deal with complex lighting and transparencies [79], and the fact that they do not rely on complex sensors. Recently, more and more rendering-based methods have been designed specifically for surface reconstruction [80, 81, 82]. They can be divided into the two categories, neural radiance fields and 3D Gaussian Splatting, and are closely related to the two corresponding, more general, principles from computer graphics, namely ray tracing and rasterization.
4.3.1 NeRF and Gaussian Splatting
NeRF encode scenes within the weights of a neural network [77] by mapping 3D locations and viewing directions with a volume density and emitted color. Although initially designed for novel view synthesis, extensions such as NeuS [83] and VolSDF [84] reparameterize the volume density using signed distance functions, allowing explicit surface extraction from radiance fields. NeRF have also been extended from object-level to entire scenes, such as Urban Radiance Fields [85], which incorporate point clouds, and Neuralangelo [86], which employs instant neural graphics primitives to improve reconstruction speed.
3D Gaussian Splatting [87] provides an efficient rendering technique based on differentiable rasterization of 3D Gaussians initialized from sparse point clouds. While primarily aimed at novel view synthesis, preliminary works like SuGaR [82] and Gaussian Frosting [88] explore efficient mesh reconstruction and real-time rendering within this framework. These methods highlight the potential of rasterization-based techniques for surface reconstruction, although they are still in the early stages.